# Leveraging data-driven weather models for improving numerical weather prediction skill through large-scale spectral nudging

### v1

@@ 4.3 Verification against radiosondes
At ECCC, model developers often rely on forecast verification against ECMWF-HRES analysis (Lang et al., 2023) to avoid the problem associated with own-analysis verification (see, e.g., Casati et al. (2008), Privé et al. (2021)). This also helps to get around the issue of inhomogeneity of the observing network. Since GC was primarily trained with the ERA5 reanalysis and then fine-tuned on the ECMWF-HRES analysis, this likely introduces some dependency on the ECMWF analysis systems. To ensure that the dataset used to verify the forecasts is completely independent from both models as well as the GDPS data assimilation system, it was opted to rely mainly on observations from the global radiosondes network. Observations of zonal wind ($U$), geopotential height ($Z$), temperature ($T$) and dewpoint depression ($T–T_{d}$) at mandatory pressure levels between 1000 and 100 hPa will therefore be used as reference for verification.
Two flavors of the verification against radiosonde observations are performed: (1) verification at full resolution, with the GDPS-CTL and GC forecasts at their native resolution, and (2) verification at low resolution –- corresponding to the effective resolution of GC – wherein both the GDPS-CTL and GC forecasts are filtered to remove the variability at scales not adequately resolved by GC. The fact that GC has a much lower effective resolution than GDPS-CTL would otherwise provide an unfair advantage to GC by reducing the so-called “double penalty” for misplacing the predicted patterns. Low-resolution verification against radiosondes also allows the study of how suitable GC forecasts are for large-scale spectral nudging in GEM.
@@
Figure 3: Verification for 120-h forecasts from GC (red) and GDPS-CTL (blue) at their full (top row: a–e) and low (bottom row: e–h) resolutions against global radiosonde observations, averaged over 60 cases in boreal winter. The error standard deviation (STDE; solid lines) and bias (mean forecast minus observation; dashed lines) are shown for (a,e) zonal wind, (b,f) geopotential heights, (c,g) temperature and (d,h) dewpoint depression. Red and blue markers denote statistically significant differences in favor of GC and GDPS-CTL, respectively. Significance is computed using the $t$-test for bias and the $F$-test for STDE. No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile. A low pass filter (Appendix A) is applied on both GC and GDPS-CTL to obtain the low-resolution fields (e–h).
@@ 4.3.1 At full resolution
Comparing the forecasts at their native resolution to the radiosonde observations reveals that for most variables and lead times, and for levels between 1000 and 200 hPa, GC forecasts are much closer to the observations than GDPS-CTL. Vertical profiles of error averages (bias) oriented as forecast minus observation and error standard deviations (STDE), calculated over the global domain for 120-h forecasts for boreal winter, highlight the typical verification differences between GC and GDPS-CTL (Fig. 3a–d). The differences between GC and GDPS-CTL in terms of biases are mixed, except for temperature during boreal winter, where clear improvements can be seen with GC. In terms of STDE, however, GC clearly shows large improvements over GDPS-CTL at almost every level, and particularly for zonal wind and dewpoint depression. Above 200 hPa, however, GC suffers from increased errors, with biases and STDE being comparable or worse than those in GDPS-CTL.
@@
with $x_{1}$ and $x_{2}$ here representing GDPS-CTL and GC, respectively. Positive values of $f_{RMSE}$ thus imply improved forecasts from GC measured by the RMSE against the radiosonde observations, while negative values represent the opposite. The forecast quality index $f_{RMSE}$ was computed every 24 h for each selected variable and level. The results for each variable were then vertically averaged for three layers: low-level (from 1000 to 850 hPa), mid-level (from 700 to 300 hPa) and high-level layer (from 250 hPa to 100 hPa).
Table 1: Latitude and longitude bounds for the 6 subdomains considered in Figures 5, 11 and 13.
@@
The region-wise variations in relative forecast performances can be examined by computing, as above, the forecast quality index $f_{RMSE}$ averaged over all variables and lead times for each vertical layer but, this time, for the 6 subdomains listed in Table 1 with their respective latitude and longitude limits. In low- and mid-level layers, GC forecasts show smaller RMSE in every subdomain in both seasons (Figs. 5a and c). In the upper-level layer, GDPS-CTL performs better during each hemisphere’s respective winter. However, GC’s poor performance during winter over the northern hemisphere is primarily originating from Europe. The largest improvements from GC in the mid- and high-level layers are found over the Tropics, which is consistent with the results of Lam et al. (2022) (see their Fig. 19).
@@ 4.3.2 At low resolution
At low resolution, the forecast improvements in terms of RMSE from GC over GDPS-CTL are significantly reduced for all variables, levels, lead times, seasons and geographical domains (Figs. 3e–h and 4). This reduction in improvements at low resolution is particularly striking in the mid-level layer over the first 72 h (compare Figs 4a and b), leading to growing forecast improvements with time from GC that saturate or peak near 120 h (Fig. 5b) compared to a relatively steady improvement at all lead times found with the full resolution (Fig. 4a). This suggests that the improvements with GC for the first 72 h at full resolution largely originate from partial smoothing of the less-accurately predicted fine scales that reduces the double penalty.
Degradations in GC, seen at full resolution over Northern Hemisphere in boreal winter (mostly coming from Europe) and Southern hemisphere in boreal summer, are enhanced at low resolution (compare Figs 5a, c, and 5b, d). There are also notable degradations in the mid- and low-level layers over Asia in summer (Fig. 6d). The average reduction factor for $f_{RMSE}$ from full- to low-resolution verification is roughly between 2 to 3 (compare the left and right columns in Fig. 5), which confirms that the lower effective resolution of GC provides a significant advantage over the GDPS-CTL. Nevertheless, large scales of GC forecasts are still, on average, significantly closer to the radiosonde observations than those from GDPS-CTL in the mid- and low-level layers. This suggests that nudging these large-scale components of key GEM state variables toward GC could result in substantial improvements over GDPS-CTL.
@@ 5.1 The concept
Spectral nudging operates by gradually directing the model-predicted atmospheric large scales towards those found in some reference data which is considered to have a more accurate representation of those scales. This approach fundamentally differs from grid nudging or indiscriminate nudging, as it only targets a predefined range of scales. Even though an NWP model may employ grid-based spatial discretizations, the nudging increments at a given model level are computed by decomposing the model predictions and the reference fields in a spectral space, retaining only the target scales. Hence, this technique is referred to as “spectral nudging”.
@@ 5.2 Implementation in GEM
With the split-adjustment-type dynamics-physics coupling in GEM, the adiabatic dynamical core first solves the prognostic dynamical equations to obtain an intermediate state of the atmosphere. In the absence of spectral nudging, this interim dynamics solution serves as the input for the physical parameterization schemes, which compute the physics-induced tendencies. These tendencies are then coupled with the dynamics solution to obtain the complete model solution for a given time step. However, when spectral nudging is applied, the large scales in the solutions for a selected set of prognostic variables from GEM dynamics are first nudged toward the reference (here, from GC). Subsequently, the nudged solutions are fed to the physics schemes to compute the physics-induced tendencies before the eventual dynamics-physics coupling, following the regular model procedure.
@@
where $F_{GEM}$ is a prognostic variable predicted by GEM dynamics, $F_{GC}$ is the corresponding prediction from GC valid at the same time, $\omega$ is the nudging relaxation factor such that $(0\leq\omega\leq 1)$, and $F_{nudge}$ is the nudged solution. The subscript $LS$ in the above equation refers to some user-defined large scales targeted by the nudging mechanism. Through simple rearrangement of the terms on the right-hand-side of Eq.4, it can be shown that the large scales in $F_{nudge}$ are indeed a weighted average of those from $F_{GEM}$ and $F_{GC}$ (with $\omega$ being the weighting parameter), whereas the small scales from $F_{GEM}$ are retained entirely in $F_{nudge}$.
@@
where $\beta(\zeta)$ defines the nudging vertical profile with respect the $\zeta$-coordinate, $\tau(t)$ represents the nudging relaxation time scale , and $\Delta t$ denotes the model time-step length (in the same unit as $\tau$). A vertically-variable $\beta$ allows for different nudging strengths at different model levels. Conversely, a time-dependent $\tau$ permits changing nudging strengths with time. In general, a large value of $\tau$ implies weak nudging, and vice versa.
To nudge a prognostic variable $F$ at any model level, it is first essential to spectrally decompose $(F_{GC}-F_{GEM})$ so that a filter can be applied in the spectral space to retain only the desirable scales. Spherical harmonics-based spectral decomposition would be the ideal choice for such a filter. However, this would require multiple transformations of model solutions between the model’s Yin-Yang grids and an intermediate global Gaussian grid, resulting in significant increase in the computational cost. As the present study is primarily of a proof-of-concept nature, in order to keep the approach simple, the spectral filter for nudging was instead chosen to be based on discrete cosine transform (DCT) (Denis et al., 2002), and the filtering is performed separately on the Yin and Yang grids. Additionally, nudging is also separately applied to the Yin and Yang grid solutions after every dynamics sub-step.
Large scales that are retained with spectral nudging are defined by users through configurable parameters. Instead of an abrupt cut-off for the target scales, which may lead to Gibbs phenomena (Sardeshmukh and Hoskins, 1984), a soft cut-off is employed within the DCT-based spectral filter. With this method, the spectral filter relies on two different cut-off scale parameters denoted by $\lambda_{LS}$ and $\lambda_{SS}$ such that scales larger than $\lambda_{LS}$ are fully retained while those smaller than $\lambda_{SS}$ are completely eliminated. A partial filtering, similar to the one utilized by Husain et al. (2014), is applied for the scales in between $\lambda_{LS}$ and $\lambda_{SS}$. Further information on the DCT-based filter and its response is provided in Appendix B.
@@ 5.3 Optimal configuration
A series of systematic sensitivity experiments was carried out to identify an optimal spectral nudging configuration for GDPS. All subsequent results presented in the following section are based on this optimal configuration which, from this point onward, will be referred to as GDPS-SN. It is important to note that spectral nudging is currently only applied within the forecasting component of GDPS, i.e., GDPS-CTL, GDPS-SN and GC are all initialized with the same analyses.
Some key aspects of this optimal setup are discussed below.
@@ 5.3.1 Nudged variables
Nudging is only applied to the $u$-$v$ components of wind and virtual temperature. While nudging specific humidity can improve the global bias of temperature and humidity in the boundary layer, the computational cost outweighs the benefits. Moreover, nudging specific humidity tends to negatively impact tropical cyclone intensity during the tested period, making it less desirable.
@@ 5.3.2 Nudging vertical profile
Although different vertical profiles, defined by $\beta(\zeta)$, have been explored, a plateau-shaped profile, as shown in Fig. 6, is found to yield the best results. This profile involves no nudging in the boundary layer (below 850 hPa) and the stratosphere (above 250 hPa). Relatively weaker performance of GC in the stratosphere – as has been mentioned earlier – is the reason for no stratospheric nudging. Within the boundary layer, nudging is avoided to prevent potential negative impacts arising from inconsistencies between the surface forcing in GDPS and GC training data (ERA5). Furthermore, GC-13 does not have sufficient vertical resolution in the boundary layer to provide a useful reference.
@@ 5.3.3 Nudging length scales
It has been already shown that the fine-scales in GC suffer from considerable smoothing, worsening at longer lead times (see Figs. 1 and 2). The implementation of spectral nudging in GEM allows for time-varying cut-off scales for nudging. However, the improvements with such an approach over a stationary cut-off scale are found to be negligible. Therefore, based on the comparison of spectra between GDPS and GC, the DCT-based filter for the optimal configuration utilized in this study is configured with $\lambda_{LS}$=2750 km and $\lambda_{SS}$=2250 km. This implies that scales larger than 2750 km are entirely retained, while those smaller than 2250 km are completely eliminated. The scales in between are subjected to partial filtering (see Fig. 20). This choice of nudging length scales corresponds to the effective resolution of GC beyond day 3 (see Figs. 1 and 2).
@@ 5.3.4 Nudging relaxation time
The nudging relaxation time, $\tau$, is analogous to the $e$-folding time as it determines the time-rate of decay in the large-scale differences between the model and the driving data in the absence of any other process. Arguably, $\tau$ is one of the most important configurable parameters for spectral nudging. However, selecting an optimal $\tau$ is not straightforward. First, it is important to note that GC inferences are only available every 6 hours and become increasingly inaccurate with longer lead times (as demonstrated by spectral coherence in Fig. 2). During model integration, when a valid GC inference is not available at certain time steps, an estimate is approximated through linear interpolation between two consecutive GC inferences around the time of interest. In this regard, Omrani et al. (2012) have argued that $\tau$ should not be less than $\tau_{a}$ which denotes the time interval between consecutive driving fields (here, GC inferences). Strong nudging with $\tau<\tau_{a}$ can lead to erroneous evolution of nudged fields. Furthermore, sensitivity experiments carried out during this study revealed considerable deterioration of model’s skill over complex terrain with $\tau<\tau_{a}$ (not shown). As a result, $\tau_{a}$ may safely be considered the lower bound for $\tau$.
It is also crucial to ascertain an acceptable upper bound for $\tau$. Intuitively, as the accuracy of reference nudging data worsens with increasing forecast lead time, it might seem prudent to employ a large $\tau$. However, sensitivity experiments have revealed that too large of a nudging relaxation time ($\tau>24$ hr) can result in excessive undesirable smoothing in the nudged fields around the cut-off scales developing at day 3 or 4. Eventually, this smoothing affects all scales by day 10 of the forecast (not shown). This increased smoothing with overly large $\tau$, is caused by the averaging effect resulting from blending two forecasts, with $\tau=48h$ maximizing the effect. Based on the insights derived from the sensitivity tests, a nudging relaxation time of 12 hr is chosen to be optimal.

### v2

@@ 4.2 Verification against radiosondes
At ECCC, model developers often rely on forecast verification against ECMWF-HRES analysis (Lang et al., 2023) to avoid the problem associated with own-analysis verification (see, e.g., Casati et al. (2008), Privé et al. (2021)). Conversely, since GC was primarily trained with the ERA5 reanalysis and then fine-tuned on the ECMWF-HRES analysis, this likely introduces some dependency on the ECMWF analysis systems. To ensure that the dataset used to verify the forecasts is as independent as possible from both models as well as the GDPS data assimilation system, it was opted to rely mainly on observations from the global radiosondes network. Observations of zonal wind ($U$), geopotential height ($Z$), temperature ($T$) and dewpoint depression ($T–T_{d}$) at mandatory pressure levels between 1000 and 100 hPa are therefore used as reference for verification.
Two flavors of verification against radiosonde observations are performed based on the results shown in the previous section: (1) verification at full resolution, with the GDPS-CTL and GC forecasts at their native resolution, and (2) verification at low resolution –- corresponding to the effective resolution of GC – wherein both the GDPS-CTL and GC forecasts are filtered to remove the variability at scales not adequately resolved by GC. The fact that GC has a much lower effective resolution than GDPS-CTL would otherwise provide an unfair advantage to GC by reducing the so-called “double penalty” for misplacing the predicted patterns (Mass et al., 2002).
@@
Figure 3: Verification for 120-hr forecasts from GC (red) and GDPS-CTL (blue) at their full (top row: a–d) and low (bottom row: e–h) resolutions against global radiosonde observations, averaged over 60 cases in boreal winter. The error standard deviation (STDE; solid lines) and bias (mean forecast minus observation; dashed lines) are shown for (a,e) zonal wind, (b,f) geopotential heights, (c,g) temperature and (d,h) dewpoint depression. Red and blue markers denote statistically significant differences in favor of GC and GDPS-CTL, respectively. Significance is computed using the $t$-test for bias and the $F$-test for STDE. No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile. A low pass filter (Appendix A) is applied on both GC and GDPS-CTL to obtain the low-resolution fields (e–h).
@@ 4.2.1 At full resolution
Comparing the forecasts at their native resolution against radiosonde observations reveals that for most variables and lead times, and for levels between 1000 and 200 hPa, GC forecasts are much closer to the observations than GDPS-CTL. Vertical profiles of error averages (bias) oriented as forecast minus observation and error standard deviations (STDE), calculated over the global domain for 120-hr forecasts for boreal winter, highlight the typical verification differences between GC and GDPS-CTL (Fig. 3a–d). The differences between GC and GDPS-CTL in terms of biases are mixed, except for temperature, where clear improvements can be seen with GC. In terms of STDE, however, GC clearly shows large improvements over GDPS-CTL at almost every level, and particularly for zonal wind and dewpoint depression. Above 200 hPa, however, GC suffers from increased errors, with biases and STDE being comparable or worse than those in GDPS-CTL.
@@
with $x_{1}$ and $x_{2}$ here representing GDPS-CTL and GC, respectively. Positive values of $f_{RMSE}$ thus imply improved forecasts from GC measured by the RMSE against the radiosonde observations, while negative values represent the opposite. The forecast quality index $f_{RMSE}$ was computed every 24 hr for each selected variable and level. The results for each variable were then vertically averaged for three layers: low-level (from 1000 to 850 hPa), mid-level (from 700 to 300 hPa) and high-level layer (from 250 hPa to 100 hPa).
Table 1: Latitude and longitude bounds for the 6 subdomains considered in Figures 5, 10 and 12.
@@
The region-wise variations in relative forecast performances can be examined by computing, as before, the forecast quality index $f_{RMSE}$ averaged over all variables and lead times for each vertical layer but, this time, for the 6 subdomains listed in Table 1 with their respective latitude and longitude limits. In low- and mid-level layers, GC forecasts show smaller RMSE in every subdomain in both seasons (Figs. 5a and c). In the high-level layer, GDPS-CTL performs better during each hemisphere’s respective winter. However, GC’s poor performance during winter over the northern hemisphere is primarily originating from Europe. The largest improvements from GC in the mid- and high-level layers are found over the Tropics, which is consistent with the results of Lam et al. (2022) (see their Fig. 19).
@@ 4.2.2 At low resolution
At low resolution, the forecast improvements in terms of RMSE from GC over GDPS-CTL are significantly reduced for all variables, levels, lead times, seasons and geographical domains (Figs. 3e–h and 4b). This reduction in improvements at low resolution is particularly striking in the mid-level layer over the first 72 hr (compare Figs 4a and b), leading to growing forecast improvements with time from GC that saturate or peak near 120 hr (Fig. 5b) compared to a relatively steady improvement at all lead times found with the full resolution (Fig. 4a). This suggests that the improvements with GC for the first 72 hr at full resolution largely originate from partial smoothing of the less-accurately predicted fine scales that reduces the double penalty.
Degradations in GC, seen at full resolution over Northern Hemisphere during boreal winter (mostly coming from Europe) and Southern hemisphere during boreal summer, are enhanced at low resolution (compare Figs 5a,c, and 5b,d). There are also notable degradations in the mid- and low-level layers over Asia in summer (Fig. 5d). The average reduction factor for $f_{RMSE}$ from full- to low-resolution verification is roughly between 2 to 3 (compare the left and right columns in Fig. 5), which confirms that the lower effective resolution of GC provides a significant advantage over the GDPS-CTL. Nevertheless, large scales of GC forecasts are still, on average, significantly closer to the radiosonde observations than those from GDPS-CTL in the mid- and low-level layers. This suggests that the GC forecasts contain useful large-scale information that could substantially improve guidance if successfully incorporated into a hybrid NWP-AI system.
@@ 5.1 The concept
Spectral nudging works by directing the model-predicted atmospheric large scales toward a more skillful reference (von Storch et al., 2000; Husain et al., 2014). This approach fundamentally differs from grid nudging or indiscriminate nudging (Liu et al., 2012), as it only targets a predefined range of scales. Even though an NWP model may employ grid-based spatial discretizations, the nudging increments at a given model level are computed by decomposing the model predictions and the reference fields in a spectral space, retaining only the target scales. Hence, this technique is referred to as “spectral nudging”.
@@ 5.2 Implementation in GEM
With the split-type dynamics-physics coupling in GEM, the adiabatic dynamical core first solves the prognostic dynamical equations to obtain an intermediate state of the atmosphere. In the absence of spectral nudging, this interim dynamics solution serves as the input for the physical parameterization schemes, which compute the physics-induced tendencies. These tendencies are then coupled with the dynamics solution to obtain the complete model solution for a given time step. However, when spectral nudging is applied, the large scales in the solutions for a selected set of prognostic variables from GEM dynamics are first nudged toward the reference (here, from GC). Subsequently, the nudged solutions are fed to the physics schemes to compute the physics-induced tendencies before the eventual dynamics-physics coupling.
@@
where $F_{GEM}$ is a prognostic variable predicted by GEM dynamics, $F_{GC}$ is the corresponding prediction from GC vertically interpolated to the $\zeta$ level valid at the same time, $\omega$ is the nudging relaxation factor such that $(0\leq\omega\leq 1)$, and $F_{nudge}$ is the nudged solution. The subscript $LS$ in the above equation refers to some user-defined large scales targeted by the nudging mechanism, and is discussed later in more detail. Through simple rearrangement of the terms on the right-hand-side of Eq.4, it can be shown that the large scales in $F_{nudge}$ are indeed a weighted average of those from $F_{GEM}$ and $F_{GC}$ (with $\omega$ being the weighting parameter), whereas the small scales from $F_{GEM}$ are retained entirely in $F_{nudge}$.
@@
where $\beta(\zeta)$ defines the nudging vertical profile with respect to the $\zeta$-coordinate, $\tau(t)$ represents the nudging relaxation time scale , and $\Delta t$ denotes the model time-step length. A vertically-variable $\beta$ allows for different nudging strengths at different model levels. Conversely, a time-dependent $\tau$ permits changing nudging strengths with time. In general, a large value of $\tau$ implies weak nudging, and vice versa.
To nudge a prognostic variable $F$ at any model level, it is first essential to spectrally decompose $(F_{GC}-F_{GEM})$ so that a filter can be applied in the spectral space to retain only the desirable scales. For global simulations, such a filter should ideally employ a spherical harmonics-based spectral decomposition. However, this would require multiple transformations of model solutions between the model’s Yin-Yang grids and an intermediate global Gaussian grid, resulting in significant increase in the computational cost. As the present study is primarily of a proof-of-concept nature, in order to keep the approach simple, the spectral filter for nudging was instead chosen to be based on discrete cosine transform (DCT) (Denis et al., 2002), and the filtering is performed separately on the Yin and Yang grids. Additionally, nudging is separately applied to the Yin and Yang grid solutions after every dynamics sub-step. It is important to note that, for the range of cut-off scales considered in this study, the response of the DCT-based filter (applied separately to the Yin-Yang grids) is found to be closely equivalent to the response of a spherical-harmonics-based filter applied to a global Gaussian field (see Appendix B).
Large scales that are retained with spectral nudging are defined through configurable parameters. Instead of an abrupt cut-off for the target scales, which may lead to Gibbs phenomena (Sardeshmukh and Hoskins, 1984), a soft cut-off is employed within the DCT-based spectral filter. With this method, the spectral filter relies on two different cut-off scale parameters denoted by $\lambda_{LS}$ and $\lambda_{SS}$ such that scales larger than $\lambda_{LS}$ are fully retained while those smaller than $\lambda_{SS}$ are completely eliminated. A partial filtering, similar to the one utilized by Husain et al. (2014), is applied for the scales in between $\lambda_{LS}$ and $\lambda_{SS}$. Further information on the DCT-based filter and its response is provided in Appendix B.
@@ 5.3 Optimal configuration
A series of systematic sensitivity experiments was carried out to identify an optimal spectral nudging configuration for GDPS. Results presented in section 6 are based on this optimal configuration which is referred to as GDPS-SN. Note that spectral nudging is currently only applied within the forecasting component of GDPS, i.e., GDPS-CTL, GDPS-SN and GC are all initialized with the same analyses.
Some key aspects of the optimal nudging configuration are discussed below.
@@ 5.3.1 Nudged variables
Nudging is only applied to the $u$-$v$ components of wind and virtual temperature. While nudging specific humidity can improve the global bias of temperature and humidity in the boundary layer, the computational cost (see section 66.6) outweighs the benefits. Moreover, sensitivity experiments have revealed a negative impact of specific humidity nudging on tropical cyclone intensity (not shown), making it less desirable.
@@ 5.3.2 Nudging vertical profile
Although different vertical profiles, defined by $\beta(\zeta)$, have been explored, a plateau-shaped profile, as shown in Fig. 6, is found to yield the best results. This profile involves no nudging in the boundary layer (below 850 hPa) and the stratosphere (above 250 hPa). Relatively weaker performance of GC in the stratosphere – as mentioned earlier – is the reason for no stratospheric nudging. Although nudging is generally avoided in the boundary layer to allow uninhibited generation of fine scales (e.g., Schubert-Frisius et al. 2017), other studies have found the potential for significant near-surface skill improvement with nudging towards an accurate reference in the boundary layer (e.g., Husain et al. 2014). However, in this study, nudging is avoided in the boundary layer for different reasons. First, the 13-level version of GC (used here) does not have sufficient vertical resolution in the boundary layer to provide a useful reference. Even in the absence of this issue, inconsistencies between the surface forcing in GDPS and GC training data (ERA5) may lead to potential negative impacts over regions with complex terrain. In addition, differences in mean boundary-layer state (moisture in particular) may result in adverse reactions from the physical parameterizations in GEM. Therefore, nudging in the boundary layer is not a viable option at this stage.
@@ 5.3.3 Nudging length scales
It was shown earlier in section 4 that the fine-scales in GC suffer from considerable smoothing, worsening at longer lead times (see Figs. 1 and 2). The implementation of spectral nudging in GEM allows for time-varying cut-off scales for nudging. However, improvements with such an approach over stationary cut-off scales are found to be negligible. Therefore, based on the comparison of spectra between GDPS and GC, the DCT-based filter for the optimal configuration utilized in this study is configured with $\lambda_{LS}$=2750 km and $\lambda_{SS}$=2250 km. This implies that scales larger than 2750 km are entirely retained, while those smaller than 2250 km are completely eliminated. The scales in between are subjected to partial filtering (see Fig. 19). This choice of nudging length scales corresponds to the effective resolution of GC beyond day 3 (see Figs. 1 and 2).
@@ 5.3.4 Nudging relaxation time
The nudging relaxation time, $\tau$, is analogous to the $e$-folding time as it determines the time-rate of decay in the large-scale differences between the model and the driving data in the absence of any other process. Arguably, $\tau$ is one of the most important configurable parameters for spectral nudging. However, selecting an optimal $\tau$ is not straightforward. First, it is important to note that GC inferences are only available every 6 hours. During model integration, when a valid GC inference is not available at certain time steps, an estimate is approximated through linear interpolation between two consecutive GC inferences around the time of interest. In this regard, Omrani et al. (2012) have argued that $\tau$ should not be less than $\tau_{a}$, which denotes the time interval between consecutive driving fields (here, GC inferences). Husain et al. (2014) showed that, with $\tau<\tau_{a}$, the impact of time interpolation error on the nudged fields may not be negligible. As a result, $\tau_{a}$ may be considered the lower bound for $\tau$.
It is also crucial to ascertain an acceptable upper bound for $\tau$. Sensitivity experiments have revealed that too large of a nudging relaxation time ($\tau>24$ hr) can result in excessive undesirable smoothing in the nudged fields around the cut-off scales developing at day 3 or 4. Eventually, this smoothing affects all scales by day 10 of the forecast (not shown). This increased smoothing with overly large $\tau$, is caused by the averaging effect resulting from blending two forecasts, with $\tau=48$ hr maximizing the effect.
Strong nudging with $\tau=\tau_{a}$ is also found to result in erroneous evolution of nudged fields emanating from inconsistencies between GEM and GC over complex terrain, as well as error originating from vertically interpolating GC to GEM levels. Therefore, based on the insights derived from the sensitivity tests, a nudging relaxation time of 12 hr is chosen to be optimal. An important point to consider in this context is that although GC inferences become increasingly inaccurate with respect to GDPS analyses at longer lead times, they still maintain higher spectral coherence (compared to GDPS-CTL) for the largest scales (see Fig. 2c,d). Therefore, a constant value of $\tau$ during the entire integration is justified.